OpenStudy (swissgirl):

Define \( \oplus \text{ on } \mathbb{R} X \mathbb{R} \) by setting \( (a,b) \oplus (c,d)=(ac-bd, ad+bc) \) Show that \( ( \mathbb{R} X \mathbb{R} , \oplus ) \) is an algebraic system

OpenStudy (swissgirl):

ohhhhh The definition is An algebraic system is an ordered pair (AO) , where A is a set, called the underlying set of the algebraic system, and O is a set, called the operator set, of finitary operations on A .

OpenStudy (swissgirl):

so I guess we gotta prove that its non empty and it has a binary operation?

OpenStudy (experimentx):

what do we have here http://planetmath.org/AlgebraicSystem.html

OpenStudy (swissgirl):

hmmmmm

OpenStudy (swissgirl):

The thing is I just checked in my book and there were 2 definitions of algebraic system which is is quite baffling

OpenStudy (swissgirl):

The first one is similar to the one stated in planet math

OpenStudy (swissgirl):

the second one imo is really the def of a group which is a kind of algebraic structure

OpenStudy (swissgirl):

like if its an algebraic system then it has the following 4 properties. commutativity, associativity, an identity element and each element has an inverse

OpenStudy (swissgirl):

ohh right and its obviously closed under that binary operation

OpenStudy (swissgirl):

haha there are way too many diff definitions in my book. i dont get it y they dont state it all at once

OpenStudy (swissgirl):

This is what someone responded but I dont think its correct -->The identity is (1, 0), since (a, b)*(1, 0)=(a(1)-b(0), a(0)+b(1))=(a, b)=(1, 0)*(a, b). But now we have a contradiction; there is no inverse for the element (0, 0), since (a, b)*(0, 0) = (0, 0) ≠ (1, 0) So, G does not form a group under *. Since the question says show then it must be an algebraic structure

OpenStudy (anonymous):

is it clear to you what this is?

OpenStudy (experimentx):

Oops!! sorry ... i totally missed. how long did OS froze?

OpenStudy (anonymous):

you want a hint?

OpenStudy (swissgirl):

yaaaaaaaa

OpenStudy (swissgirl):

Like I am not sure exactly what an algebraic structure is

OpenStudy (anonymous):

here is my hint: let \(z_1=a+bi,z_2=c+di\) be two complex numbers. what is \(z_1\times z_2\)?

OpenStudy (anonymous):

i just read your post above. one does not expect to find an inverse for the zero element, but you should be able to find an inverse for anything else

OpenStudy (experimentx):

An algebraic system is an ordered pair (A,O) , where A is a set, called the underlying set of the algebraic system, and O is a set, called the operator set, of finitary operations on A .

OpenStudy (experimentx):

I bet RxR is A

OpenStudy (experimentx):

for O, we have ⊕

OpenStudy (swissgirl):

right

OpenStudy (experimentx):

it isn't difficult to prove that it's closed and commutative.

OpenStudy (experimentx):

(1, 0) is the identity element inverse ... yep here is the trouble (0,0) does not have inverse.

OpenStudy (swissgirl):

but what was satellite's hint

OpenStudy (swissgirl):

How does that come into play?

OpenStudy (experimentx):

let's try this Q before is (R, *) a group?

OpenStudy (swissgirl):

yes

OpenStudy (anonymous):

did you multiply \((a+bi)(c+di)\) ?

OpenStudy (experimentx):

i guess yes http://in.answers.yahoo.com/question/index?qid=20091104013411AA3mtNV from planet math: "A prototypical example of an algebraic system is a group, which consists of the underlying set G , and a set O consisting of three operators: a constant e called the multiplicative identity, a unary operator called the multiplicative inverse, and a binary operator called the multiplication."

OpenStudy (anonymous):

\((\mathbb{R},\times)\) is not a group because zero has no inverse. you do not expect 0 (the identity under addition) to have a multiplicative inverse

OpenStudy (swissgirl):

So then what abt this example?

OpenStudy (experimentx):

so what do we say here? on this particular example ... i'm not sure either ...algebra has never been my thing.

OpenStudy (anonymous):

the zero element here, namely \((0,0)\) has no inverse, you would not expect it to my hint was to get you to see that this is the beginning of the construction of complex numbers from the real numbers

OpenStudy (experimentx):

perhaps not http://en.wikipedia.org/wiki/Group_(mathematics)#Definition says "each element must have multiplicative inverse"

OpenStudy (anonymous):

you start by defining a set \(\mathbb{R}\times \mathbb{R}\) @experimentX that is for a group, this is not a group

OpenStudy (experimentx):

why ... does algebraic system require further criteria?

OpenStudy (swissgirl):

Like what is an algebraic structure what are the properties of an algebraic structure?

OpenStudy (anonymous):

you define addition for \(\mathbb{R}\times \mathbb{R}\) to be \((a,b)+(c,d)=(a+c,b+d)\) and multiplication to be defined as above namely \((a,b), (c,d)=(ac-bc,ad+bc)\) then you show \((0,1)\times (0,1)=(-1,0)\) you can then show that this is a field, and finally you define \((0,1)=i\) and rewrite the ordered pair \((a,b)\) as \(a+bi\)

OpenStudy (anonymous):

@swissgirl work directly from the definition, whatever definition you have i think there is not really much to show you are not expected to show that this is a group (because it isn't) only that you have a set and a binary operation

OpenStudy (anonymous):

ok here is what you wrote as the definition The definition is An algebraic system is an ordered pair (AO) , where A is a set, called the underlying set of the algebraic system, and O is a set, called the operator set, of finitary operations on A so in this example \[A=\mathbb{R}\times \mathbb{R}\] and \(O=\oplus\)

OpenStudy (anonymous):

since \(\mathbb{R}\times \mathbb{R}\) is a set, and since \(\oplus\) is a binary operation (takes two elements of the set and send it to one element in the set) then by definition this is an algebraic structure there is really no more to be done

OpenStudy (swissgirl):

but what were u showing earlier

OpenStudy (experimentx):

but what's the difference between algebraic structure and algebraic system?

OpenStudy (swissgirl):

I think its the same thing

OpenStudy (experimentx):

I googled and got this pdf with Galois stuff ...

OpenStudy (anonymous):

me? i was showing that this is how on constructs complex numbers out of real numbers, by starting with ordered pairs but that is not what the question asks, so you can ignore that part